Margo borrows $1300, agreeing to pay it back with 4% annual interest after 13 months. How much interest will she pay?
Round your answer to the nearest cent, if necessary.

Answer by stanbon(75887)

Margo borrows $1300, agreeing to pay it back with 4% annual interest after 13 months. How much interest will she pay?
Round your answer to the nearest cent, if necessary.
------
I = P*r*t
Int = 1300*0.04*(13/12) = $56.33
--------------
Cheers,
Stan H.
---------